The product of two consecutive odd integers is 575 what are the integers.
x= 1st number
x+2=2nd number

Answer by checkley75(3666)

x(x+2)=575
x^2+2x-575=0
(x+25)(x-23)=0
x-23=0
x=23 for the smaller integer.
23+2=25 for the larger integer.
Proof:
23*25=575
575=575